# From Coarse to Fine-Grained Emotion Annotation: An Immediate Recall Paradigm with Validation through Physiological Evidence and Recognition Performance

### v1

@@ Abstract
Traditional video-induced emotion physiological datasets often use whole-trial annotation, assigning a single emotion label to all data collected during an entire trial. This coarse-grained annotation approach misaligns with the dynamic and temporally localized nature of emotional responses as they unfold with video narratives, introducing label noise that limits emotion recognition algorithm evaluation and performance. To solve the label noise problem caused by coarse-grained annotation, we propose a fine-grained annotation method through an immediate recall paradigm. This paradigm integrates an immediate video replay phase after the initial stimulus viewing, allowing participants to precisely mark the onset timestamp, emotion label, and intensity based on their immediate recall. We validate this paradigm through physiological evidence and recognition performance. Physiological validation of multimodal signals within participant-marked windows revealed rhythm-specific EEG patterns and arousal-dependent GSR responses—with SCRs appearing in 91% of high-arousal versus 6% of low-arousal emotion windows. These objective physiological data changes strongly aligned with subjective annotations, confirming annotation precision. For recognition performance, classification experiments showed that models trained on fine-grained annotations achieved 9.7% higher accuracy than traditional whole-trial labeling, despite using less data. This work not only addresses label noise through fine-grained annotation but also demonstrates that annotation precision outweighs data scale in determining emotion recognition performance.
Index Terms:
Emotion Annotation, Label Noise, Multimodal Emotion Dataset, Physiological Validation, Emotion Recognition.
@@ I Introduction
Emotion is a subjective experience elicited by external stimuli [1], typically accompanied by physiological responses and behavioral tendencies. Emotions play a crucial role in individual survival, social interaction [2], and decision-making processes [3]. Two primary theoretical frameworks describe emotions: discrete and dimensional models. The discrete model, exemplified by Ekman’s seminal work [4], identifies six basic emotions—happiness, sadness, fear, anger, surprise, and disgust—each linked to specific neural circuits. For instance, fear demonstrates strong associations with amygdala activation. In contrast, the dimensional model, represented by Russell’s Circumplex Model [5], characterizes emotions along two dimensions: valence and arousal. These complementary frameworks establish a robust theoretical foundation for emotion research.
@@
To explore neural mechanisms and develop recognition methods for emotions, researchers have designed various experimental paradigms to simulate and capture human emotional responses. These experiments typically employ music [6], images [7], or videos [8] as stimuli while collecting participants’ physiological and behavioral signals. Among these modalities, video stimuli demonstrate superior performance in eliciting complex emotions due to their dynamic nature and multisensory input, thereby offering higher ecological validity compared to static images or single-modality audio stimuli [9]. Data collection encompasses both physiological signals (e.g., electroencephalogram (EEG), electrocardiogram (ECG)) [10, 11] and non-physiological signals (e.g., facial expressions) [12], providing comprehensive documentation of emotional manifestations. Physiological signals offer high temporal resolution and objectivity, enabling precise capture of neural and physiological changes associated with emotions [13]. Non-physiological signals provide behavioral cues that complement physiological measures, further deepening our multidimensional understanding of emotional mechanisms [14].
@@
Despite continuous developments in experimental paradigms for emotion research, existing methods face significant challenges in emotion annotation accuracy. Datasets such as SEED [15] and DEAP [16] employ coarse-grained annotation strategies, assigning single emotion labels to entire multi-minute trials. While this approach reduces annotation time, it overlooks the temporal instability and discrete nature of emotions, resulting in substantial label noise. To address this limitation, Zhang et al. [17] proposed an emotion localization task for the SEED dataset. They segmented each video stimulus into multiple intervals and recruited 20 annotators to evaluate whether each segment triggered the target emotion. Segments receiving majority agreement were labeled as ground truth, thereby reducing noise by focusing on highly relevant emotional segments. Building upon this foundation [17], Zhang et al. [18] introduced TAS-Net, an unsupervised time-aware sampling network based on deep reinforcement learning. This approach automatically detects key emotional segments in EEG signals while reducing labeling noise from traditional fixed-label methods. The method’s accuracy was validated by comparing detected segments with manually re-annotated segments. Similarly, Jiang et al. proposed REMONET [19], which mitigates sentiment label noise through a multi-regularization co-learning framework that coordinates different self-supervised data views to prevent model overfitting to noisy labels. However, these post-hoc approaches—whether through manual re-annotation or deep learning techniques—do not fundamentally resolve the inherent ambiguity in mapping physiological data to subjective emotion labels [20].
In recent efforts to address annotation granularity, researchers have developed multimodal datasets such as SEED-VII [21] and MGEED [22], aiming to achieve finer-grained emotional labeling. SEED-VII represents a significant advancement by incorporating both trial-by-trial intensity ratings and post-session continuous annotation. Specifically, participants provide immediate intensity feedback after each trial, then retrospectively assign frame-level emotional intensities while reviewing all videos at the session’s conclusion. This dual-annotation approach attempts to capture both immediate impressions and detailed temporal dynamics. Similarly, MGEED employs a comprehensive frame-by-frame retrospective protocol, where participants meticulously annotate their emotional states for each video frame after viewing all experimental stimuli. While these methodologies achieve fine-grained temporal annotation, both datasets fundamentally rely on delayed recall mechanisms. This “watch first, annotate later” paradigm, despite its temporal granularity, introduces significant cognitive load [23], as participants must accurately remember and differentiate emotional experiences across multiple stimuli viewed over extended periods. Moreover, inherent memory bias [24] in delayed reporting may lead to temporal displacement of emotional events or conflation of distinct emotional episodes, potentially compromising annotation reliability.
To achieve fine-grained emotion annotation while mitigating memory bias, we propose an immediate recall annotation paradigm and establish the Fine-grained Immediate Recall-based Multimodal Emotion Dataset (FIRMED). This approach introduces three key methodological advancements:
1) Precise Temporal Localization via Immediate Recall: To achieve precise temporal localization of emotional events within each trial, our paradigm incorporates an immediate replay phase following stimulus presentation. During post-stimulus replay, participants retrospectively identify and mark the discrete time point $t_{\text{event}}$ at which they recall first experiencing the target emotion during initial viewing. A fixed-duration window (4 seconds) centered at $t_{\text{event}}$ is then extracted for analysis, associating data within this window with the reported emotion label and intensity. This recall-based discrete annotation during immediate replay aims to overcome temporal ambiguity in whole-trial labeling while reducing memory decay biases associated with delayed annotation protocols.
2) Expanded Multimodal Data for Enhanced Reliability: To improve annotation reliability and objectivity, we expanded the range of collected physiological signals to include EEG, ECG, galvanic skin response (GSR), and photoplethysmogram (PPG) data. Additionally, we captured participants’ facial expressions via camera during stimulus viewing, providing behavioral data on emotional responses. This multimodal collection strategy enriches data variety and enables cross-validation of subjective annotations using various physiological data.
3) Ecologically Valid Stimuli: We extensively utilize short videos from social media platforms as stimuli, encompassing authentic scenes including daily life recordings and news reports. Compared to traditional movie clips or music videos, these stimuli more accurately reflect natural emotional experiences, offering enhanced ecological validity. Participant feedback regarding stimulus effectiveness further supports this choice.
@@
To validate this paradigm’s effectiveness, we conducted systematic physiological and computational analyses. Physiologically, by analyzing multimodal signals within precisely annotated windows, we found strong evidence of both central [25] and autonomic nervous system engagement [26]. This included discovering emotion-specific spectral patterns in EEG signals and strong temporal coupling between discrete annotations and objective autonomic nervous system activation markers, exemplified by specific skin conductance response (SCR) patterns reliably occurring for high-arousal emotions [27]. Computationally, emotion classification experiments compared the effectiveness of the two labeling strategies. Results demonstrate that standard classifiers trained using precise labels from FIRMED significantly outperform identical models trained using traditional whole-trial labels applied to corresponding full trial segments from the same data, thereby quantifying the positive impact of reduced label noise on recognition performance.
@@ II Related Work
II-A Evolution of Physiological Signal-Based Emotion Datasets
Video-based emotion elicitation has established itself as the predominant stimulus modality for inducing emotional responses in controlled experimental environments, driving the proliferation of multimodal datasets across the affective computing domain [28]. These datasets systematically collect multiple physiological signals to provide comprehensive documentation of participants’ emotional states [29].
Early and highly influential datasets, including DEAP [16] and the foundational SEED series [15, 30, 31], established the methodological foundation for physiological emotion analysis through rich signal acquisition protocols. These seminal works adopted a standardized experimental paradigm involving presentation of emotionally evocative stimuli—typically music videos or film clips—followed by collection of self-reported emotional assessments. However, these datasets universally employ whole-trial labeling, wherein single emotion labels are assigned to complete multi-minute stimulus segments.
This coarse-grained annotation strategy ignored the inherently dynamic nature of video stimuli. Emotional responses are characteristically triggered by specific narrative elements—including pivotal plot developments, climactic sequences, or sudden scene transitions—rather than being uniformly sustained throughout extended video segments [32]. Consequently, this labeling granularity introduces label noise [19], as temporal segments that fail to elicit the target emotion become erroneously associated with that emotional state.
In direct response to these temporal granularity limitations, recent dataset development efforts have pursued increasingly sophisticated annotation methodologies. The SEED-VII dataset [21] introduced a delayed continuous rating paradigm representing a significant methodological advancement. In this approach, participants retrospectively review all trial videos upon session completion and provide continuous intensity annotations. Similarly, the MGEED database [22] implemented a delayed frame-by-frame self-assessment protocol, requiring participants to annotate their emotional states retrospectively following complete viewing session.
To systematically characterize these dataset developments, Table I provides a comprehensive comparison of representative emotion datasets. The analysis reveals a clear evolutionary trajectory from coarse whole-trial labeling toward increasingly fine-grained temporal annotations.
@@ II-A Evolution of Physiological Signal-Based Emotion Datasets
EEG (62-ch), EOG 6 Discrete emotions + Neutral Continuous
@@
Abbreviations: VAD = Valence-Arousal-Dominance; RSP = Respiration; BVP = Blood Volume Pulse; Temp. = Temperature; A-V = Audio-Visual
II-B Challenges of Emotion Annotation
The systematic adoption of whole-trial labeling across early multimodal emotion datasets [16, 15, 34, 33, 8, 35]. This annotation approach suffers from temporal imprecision: emotions constitute dynamic phenomena that evolve synchronously with video narratives [32], exhibiting rapid fluctuations and transitions rather than maintaining static states across extended durations. The assignment of uniform labels to multi-minute segments obscures emotional transitions and fluctuations [36], causing label noise to be structurally embedded in the dataset.
While recent datasets such as SEED-VII and MGEED have achieved significant improvements in labeling granularity through delayed recall paradigms. These approaches require participants to complete entire experimental sessions—encompassing numerous stimuli viewed over extended periods (typically 30 minutes to over 1 hour)—before retrospectively annotating their emotional states across all preceding trials. This delayed annotation strategy, despite achieving fine temporal granularity, it also increases the risk of memory bias, which may reduce the reliability of annotations. Cognitive psychology research [37] demonstrates that episodic memory for specific emotional events and their nuanced temporal characteristics becomes increasingly unreliable when recall is required after extended delays involving multiple distinct experiences.
The persistence of these limitations—temporal imprecision in whole-trial labeling and memory bias in delayed fine-grained annotation—underscores the urgent need for innovative annotation paradigms capable of achieving fine temporal granularity while maintaining annotation reliability.
@@ III-A Stimuli
To enhance ecological validity and elicit emotional responses more representative of real-world experiences, we primarily selected short videos sourced from social media platforms (e.g., Douyin, Bilibili, YouTube) as stimuli, each lasting approximately one minute. These videos were often user-generated or edited by creators. The content primarily included daily life recordings and news reports, and the videos typically contained multiple segments potentially capable of evoking specific emotions within the short timeframe.
The stimulus selection involved a rigorous screening process. First, a large pool of candidate videos was initially collected and downloaded from various online platforms by two members of the research team. Subsequently, these candidates were strictly evaluated by a panel of three independent raters who are psychology graduate students trained in emotion recognition. The selection criterion was: only videos for which all three raters unanimously agreed could effectively elicit one specific target basic emotion (Happiness, Sadness, Anger, Fear, Disgust, or Surprise) were ultimately chosen. Videos failing to reach unanimous agreement on a single target emotion or eliciting mixed emotions were excluded during this phase.
Through this rigorous screening process, we obtained the final stimulus set containing 90 selected videos. This set was balanced across the six basic emotions and these 90 videos were evenly distributed across 3 experimental sessions. Each session included 30 trials, with each trial featuring one video specifically selected to elicit one of the six target basic emotions, resulting in 5 distinct videos per emotion category per session. This stimulus selection strategy, prioritizing ecological validity and confirming elicitation of specific target emotions via consensus rating, was corroborated by informal post-experimental feedback from participants; they generally reported that these real-world clips were more engaging and significantly more effective in eliciting their emotional resonance compared to more traditional laboratory stimuli such as movie clips or music videos.

### v2

@@ Abstract
Traditional video-induced emotion physiological datasets typically employ whole-trial annotation, assigning a single emotion label to all data collected throughout the experiment. This coarse-grained annotation method is inconsistent with the dynamic and temporally localized nature of emotional responses within video narratives, introducing label noise. To address the label noise problem caused by coarse-grained annotation, inspired by the Peak-End Rule, we propose a fine-grained annotation method based on an immediate recall paradigm. This paradigm incorporates a video replay phase immediately after the initial stimulus viewing, allowing participants to precisely annotate the timestamp, emotion label, and intensity of emotional outbursts based on immediate recall. We validate this paradigm through physiological evidence and recognition performance. Physiological validation of multimodal signals within participant-marked windows revealed rhythm-specific EEG patterns; for instance, fear exhibits the most widespread $\alpha$ desynchronization, while happiness shows significant $\gamma$ and $\beta$ enhancement. Meanwhile, arousal-dependent GSR responses were observed—with SCRs appearing in 79.8% of high-arousal versus 12.7% of low-arousal emotion windows. These objective physiological data changes strongly aligned with subjective annotations, confirming annotation precision. For recognition performance, classification experiments showed that models trained on fine-grained annotations achieved 9.1% higher accuracy than traditional whole-trial labeling, despite using less data. This work not only addresses label noise through fine-grained annotation but also demonstrates that annotation precision outweighs data scale in determining emotion recognition performance.
@@ I Introduction
Emotion is a subjective experience elicited by external stimuli [frijda2005emotion], typically accompanied by physiological responses and behavioral tendencies. Emotions play a crucial role in individual survival[rolls2014emotion]. Two primary theoretical frameworks describe emotions: discrete and dimensional models. The discrete model, exemplified by Ekman’s seminal work [ekman1999basic], identifies six basic emotions—happiness, sadness, fear, anger, surprise, and disgust—each linked to specific neural circuits. For instance, fear demonstrates strong associations with amygdala activation. In contrast, the dimensional model, represented by Russell’s Circumplex Model [russell1980circumplex], characterizes emotions along two dimensions: valence and arousal. These complementary frameworks establish a robust theoretical foundation for emotion research.
@@
To explore neural mechanisms and develop recognition methods for emotions, researchers have designed various experimental paradigms to simulate and capture human emotional responses. These experiments typically employ music [eerola2012review], images [junghofer2001fleeting], or videos [soleymani2011multimodal] as stimuli while collecting participants’ physiological and behavioral signals. Among these modalities, video stimuli demonstrate superior performance in eliciting complex emotions due to their dynamic nature and multisensory input, thereby offering higher ecological validity compared to static images or single-modality audio stimuli [siedlecka2019experimental]. Data collection encompasses both physiological signals (e.g., electroencephalogram (EEG), electrocardiogram (ECG)) [alarcao2017emotions, hsu2017automatic, shen2025physiological] and non-physiological signals (e.g., facial expressions) [liu2017facial], providing comprehensive documentation of emotional manifestations. Physiological signals offer high temporal resolution and objectivity, enabling precise capture of neural and physiological changes associated with emotions [shu2018review]. Non-physiological signals provide behavioral cues that complement physiological measures, further deepening our multidimensional understanding of emotional mechanisms [pillalamarri2025review].
@@
Despite continuous developments in experimental paradigms for emotion research, existing methods face significant challenges in emotion annotation accuracy. Datasets such as SEED [zheng2015investigating] and DEAP [koelstra2011deap] employ coarse-grained annotation strategies, assigning single emotion labels to entire multi-minute trials. While this approach reduces annotation time, it overlooks the temporal instability and discrete nature of emotions, resulting in substantial label noise. To address this limitation, Zhang et al. [zhang2022eeg] proposed an emotion localization task for the SEED dataset. They segmented each video stimulus into multiple intervals and recruited 20 annotators to evaluate whether each segment triggered the target emotion. Segments receiving majority agreement were labeled as ground truth, thereby reducing noise by focusing on highly relevant emotional segments. Building upon this foundation [zhang2022eeg], Zhang et al. [zhang2023unsupervised] introduced TAS-Net, an unsupervised time-aware sampling network based on deep reinforcement learning. This approach automatically detects key emotional segments in EEG signals while reducing labeling noise from traditional fixed-label methods. The method’s accuracy was validated by comparing detected segments with manually re-annotated segments. Similarly, Jiang et al. proposed REMONET [jiang2024remonet], which mitigates sentiment label noise through a multi-regularization co-learning framework that coordinates different self-supervised data views to prevent model overfitting to noisy labels. However, these post-hoc approaches—whether through manual re-annotation or deep learning techniques—do not fundamentally resolve the inherent ambiguity in mapping physiological data to subjective emotion labels [winter1997individual].
Recent efforts like SEED-VII [jiang2024seed] and MGEED [wang2023mgeed] improved temporal granularity using delayed, continuous, or frame-by-frame annotations. However, these “watch first, annotate later” paradigms fundamentally rely on delayed recall over extended periods, incurring vulnerability to memory bias [thomas1990memory, colombo2020affect]. This limitation stems from the human mechanism of emotional memory. Research by Daniel Kahneman, particularly the Peak-End Rule, posits that retrospective memory disproportionately retains ’snapshots’ of peak and end moments, rather than a continuous, faithful recording [kahneman1993more, fredrickson1993duration, horwitz2024review]. Consequently, demanding continuous retrospective annotation forces reconstruction from an inherently biased memory system, potentially more insidious form of label noise despite the fine granularity.
To overcome the twin challenges of temporal imprecision in coarse-grained methods and memory bias in delayed fine-grained methods, this paper proposes a novel immediate recall annotation paradigm specifically designed to capture high-fidelity emotional moments. We introduce the Fine-grained Immediate Recall-based Multimodal Emotion Dataset (FIRMED), created using this paradigm. By incorporating an immediate replay phase after each stimulus, our method empowers participants to precisely identify and mark the discrete timestamp ($t_{\text{event}}$) of subjectively experienced emotional peaks based on fresh recall. This directly targets the most robustly encoded ’snapshots’ suggested by the Peak-End Rule, aiming to minimize label noise at the source.
We further bolster the quality and utility of FIRMED through:
1) Expanded Multimodal Recordings: Synchronized EEG, ECG, Galvanic Skin Response (GSR), photoplethysmogram (PPG), and facial video enable robust cross-validation between subjective labels and objective physiological responses.
2) Ecologically Valid Stimuli: Traditional datasets using film clips may introduce a cognitive load of narrative comprehension [kimura2019emotional, guo2022emotional]. In contrast, we selected short videos from social media as our stimuli. Most of these videos, consisting of news reports or daily life recordings, can be understood without imposing significant extra cognitive load.
@@
To validate this paradigm’s effectiveness, we conducted systematic physiological and computational analyses. Physiologically, by analyzing multimodal signals within precisely annotated windows, we found strong evidence of both central [hagemann2003central] and autonomic nervous system engagement [kreibig2010autonomic]. This included discovering emotion-specific spectral patterns in EEG signals and strong temporal coupling between discrete annotations and objective autonomic nervous system activation markers, exemplified by specific skin conductance response (SCR) patterns reliably occurring for high-arousal emotions [xia2017salience]. Computationally, emotion classification experiments compared the effectiveness of the two labeling strategies. Results demonstrate that standard classifiers trained using precise labels from FIRMED significantly outperform identical models trained using traditional whole-trial labels applied to corresponding full trial segments from the same data, thereby quantifying the positive impact of reduced label noise on recognition performance.
@@ II Related Work
II-A Evolution of video-induced Physiological Signal-Based Emotion Datasets
Video-induced emotion elicitation has established itself as the predominant stimulus modality for inducing emotional responses in controlled experimental environments [pei2025identifying], driving the proliferation of multimodal datasets across the affective computing domain [li2022eeg]. These datasets systematically collect multiple physiological signals to provide comprehensive documentation of participants’ emotional states [wang2023unlocking].
Early and highly influential datasets, including DEAP [koelstra2011deap] and the foundational SEED series [zheng2015investigating, zheng2018emotionmeter, liu2021comparing], established the methodological foundation for physiological emotion analysis through rich signal acquisition protocols. These seminal works adopted a standardized experimental paradigm involving presentation of emotionally evocative stimuli—typically music videos or film clips—followed by collection of self-reported emotional assessments. However, these datasets universally employ whole-trial labeling, wherein single emotion labels are assigned to complete multi-minute stimulus segments. This coarse-grained annotation strategy ignored the inherently dynamic nature of video stimuli. Emotional responses are characteristically triggered by specific narrative elements—including pivotal plot developments, climactic sequences, or sudden scene transitions—rather than being uniformly sustained throughout extended video segments [hakim2013computational]. Consequently, this labeling granularity introduces label noise [jiang2024remonet], as temporal segments that fail to elicit the target emotion become erroneously associated with that emotional state.
In direct response to these temporal granularity limitations, recent dataset development efforts have pursued increasingly sophisticated annotation methodologies. Datasets such as SEED-VII [jiang2024seed] and MGEED [wang2023mgeed] represent significant advancements, employing delayed recall paradigms like continuous ratings or frame-by-frame assessments performed after the entire experimental session to achieve fine-grained labeling. While these approaches improve temporal resolution compared to whole-trial methods, they introduce a critical challenge: memory bias. Research of the Peak-End Rule, suggests that our retrospective memory of an affective experience is not a continuous, faithful recording but is disproportionately dominated by its most intense moment (the “peak”) and its final moment (the “end”) [kahneman1993more, fredrickson1993duration]. Therefore, asking participants to retrospectively draw a continuous intensity curve forces them to reconstruct a timeline their memory has not accurately stored, potentially introducing label noise despite the fine granularity.
To systematically characterize these dataset developments, Table I provides a comprehensive comparison of representative emotion datasets. The analysis reveals a clear evolutionary trajectory from coarse whole-trial labeling toward increasingly fine-grained temporal annotations. It can be seen that exploring more effective emotion labeling paradigms is still a key issue worthy of research in affective computing.
@@ II-A Evolution of video-induced Physiological Signal-Based Emotion Datasets
EEG (62-ch), EOG 6 Discrete emotions + Neutral Emotion category and intensity labels for each 4-s time windows
@@
Legend: VAD = Valence-Arousal-Dominance; RSP = Respiration; BVP = Blood Volume Pulse; Temp. = Temperature; A-V = Audio-Visual
II-B Emotion Recognition Based on Physiological Signals
Alongside the evolution of datasets, the algorithms for classifying emotions from physiological signals have also advanced significantly. This development has progressed along two primary tracks: single-modality analysis, which focuses on rich signals like EEG, and multimodal fusion, which integrates information from multiple physiological sources to achieve a more comprehensive understanding of the emotional state[chen2023multi].
In the single-modality (primarily EEG-based) track, the field has transitioned from traditional machine learning methods like Support Vector Machines (SVM) [wang2011eeg] and k-Nearest Neighbors (k-NN) [li2018emotion]—which often rely on hand-crafted features like Differential Entropy (DE) [duan2013differential]—to end-to-end deep learning. Architectures such as EEGNet [lawhern2018eegnet] capture spatio-temporal features, while Recurrent Neural Networks (RNNs) like Long Short-Term Memory (LSTM) [graves2012long] model the temporal dependencies. More recently, Graph Convolutional Neural Networks (GCNs) have shown strong performance by modeling the brain’s functional connectivity as a dynamic graph [song2018eeg, tang2024multi].
However, emotions are complex, multi-system phenomena, involving both the Central Nervous System (CNS) [shen2025ua], captured by EEG, and the Autonomic Nervous System (ANS) [shen2025physiological], reflected in peripheral signals [kreibig2010autonomic]. Relying on EEG alone may miss crucial aspects of the emotional response, such as changes in physiological arousal and sympathetic nervous system. Consequently, a critical and dominant trend in the field is multimodal emotion recognition. This approach seeks to achieve more robust and accurate classifications by fusing complementary data streams. Common strategies involve combining EEG with peripheral signals like GSR [goshvarpour2017accurate], ECG [wang2023novel], and Eye movement [lu2015combining, gong2024ciabl]. As this fusion becomes more complex, advanced architectures are being explored. Li et al. proposed the UAGCFNet [li2025uncertainty], it employs an Uncertainty-Aware Graph Convolutional Network (UAGCN) based on probabilistic modeling to estimate and quantify the uncertainty within each modality’s graph representation. Furthermore, a Transitive Contrastive Fusion (TCF) module, combining Criss-Cross Attention (CCA) and Transitive Contrastive Learning (TCL), is introduced to mitigate inter-modal uncertainty and eliminate unimodal bias during the fusion process. To explicitly tackle the spatial heterogeneity intrinsic to EEG signals and the temporal heterogeneity across diverse physiological modalities , Ma et al. developed the MSTM-Net [ma2025multilevel]. The network utilizes a Holistic Spatial Feature Aggregator (HSFA) to capture both local hemispheric differences and global spatial patterns from EEG data through multiscale feature learning. Additionally, a Hierarchical Temporal Fusion Module (HTFM) progressively integrates auxiliary physiological signals with the primary EEG modality using multiple attention-based fusion levels to address temporal disparities effectively.
@@ III-A Stimuli
To enhance ecological validity and elicit emotional responses more representative of real-world experiences, we primarily selected short videos sourced from social media platforms (e.g., Douyin, Bilibili, YouTube) as stimuli, each lasting approximately one minute. These videos were often user-generated or edited by creators. Unlike traditional film clips, our stimuli consisted primarily of daily life recordings and news reports, which are intuitive and do not require complex plot comprehension. The videos typically contained multiple segments potentially capable of evoking specific emotions within the short timeframe.
The stimulus selection involved a rigorous screening process. First, a large pool of candidate videos was initially collected and downloaded from various online platforms by two members of the research team. Subsequently, these candidates were strictly evaluated by a panel of three independent raters who are psychology graduate students trained in emotion recognition. The selection was guided by two primary criteria. First, only videos for which all three raters unanimously agreed could effectively elicit one specific target basic emotion (Happiness, Sadness, Anger, Fear, Disgust, or Surprise) were ultimately chosen.Second, to balance ecological validity with experimental control, videos containing excessive potential confounders (e.g., distracting text overlays or overly complex editing) were explicitly deprioritized during this phase. Videos failing to reach unanimous agreement on a single target emotion or eliciting mixed emotions were excluded during this phase.
Through this rigorous screening process, we obtained the final stimulus set containing 90 selected videos. This set was balanced across the six basic emotions and these 90 videos were evenly distributed across 3 experimental sessions. Each session included 30 trials, with each trial featuring one video specifically selected to elicit one of the six target basic emotions, resulting in 5 distinct videos per emotion category per session. This stimulus selection strategy, prioritizing ecological validity and confirming elicitation of specific target emotions via consensus rating, was corroborated by informal post-experimental feedback from participants; they generally reported that these real-world clips were more engaging and significantly more effective in eliciting their emotional resonance compared to more traditional laboratory stimuli such as movie clips or music videos, a finding we attribute to the reduced cognitive load and more direct emotional impact of our chosen stimuli.